HUBBLE SPACE TELESCOPE DAILY REPORT #1779 PERIOD COVERED: 0000Z (UTC) 12/13/96 - 0000Z (UTC) 12/16/96 Daily Status Report as of 351/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Two Sets of WF/PC-2 6253 (Pure Parallel Archive Program) The WF/PC-2 was used to execute a pure parallel program taking = ultraviolet observations running concurrently with prime science = observations. The intent is to search for extra-galactic sources with = flux excesses at 3000 A0 and galactic sources with excesses at 1600 A0. = The observations were completed as planned, and no problems were = reported. 1.2 Completed WF/PC-2 6362 (Supernova Host Galaxy) The WF/PC-2 was used to make observations of the supernova host = galaxy GAL-122852. The observations were executed as scheduled, and no = anomalies were noted. 1.3 Completed WF/PC-2 6351 (Resolving SDB Binary Systems) The WF/PC-2 was used to make high resolution images of the sdB = class, binary system STAR-22817 in order to try and resolve its two = individual stars. sEvidence has accumulated that a large fraction (if not = all) of the sdB class stars are found in binary systems. The = observations were executed as scheduled, and no anomalies were noted. 1.4 Completed FOS/BL 6465 (The UV Upturn of Super-Metal-Rich Globular = Clusters) The Faint Object Spectrograph (Blue Side) was used to make = ultraviolet spectrographic observations of the super-metal-rich globular = cluster NGC-224-G17 which is located within the Andromeda Galaxy (aka, = M-31) in order to observe the spectral energy distribution of this = cluster. The objective is to determine whether the strong ultraviolet = fluxes seen in the galaxy can also be seen in this cluster. The observati= ons were executed as scheduled, and no problems were reported 1.5 Completed WF/PC-2 7203 (Cosmology with the Medium Deep Survey - = Short Orbit Cases) The WF/PC-2 was used to take short (about 20 minute) exposures, to = be taken in parallel with the prime science observations, using the I = filter. These exposures are part of the medium-deep survey, which = emphasizes exposures that will be useful in obtaining cosmological = parameters. The observations were completed as planned, and no problems w= ere reported. 1.6 Completed FOS/BL 6911 (FOS Cycle 6: Discriminator Test) The Faint Object Spectrograph (Blue Side) was used to execute a = discriminator calibration test. Because both the noise and gain are = known to be temperature sensitive, it is likely that some fraction of the = channels will experience some change in their optical discriminator = settings on orbit. An internal flat calibration image was also taken. = The observations were completed as planned, and no anomalies were = reported. 1.7 Completed GHRS/2 6455 (Metal Diffusion and Radiative Levitation in = Hot Helium Rich White Dwarfs) The High Resolution Spectrograph (Side Two) was used to make = spectrographic observations of the hot, helium-rich, white dwarf star = HZ21 in order to study its metal diffusion and radiative levitation. The = data will be used to search for weak photospheric metal absorption lines = which were too weak to be observed at optical wavelengths. The = observations were executed as planned, and no anomalies were reported. 1.8 Completed FOS/BL 6799 (The BAL Region Covering Factor in a Sample = of IRAS-Selected QSOs) The Faint Object Spectrograph (Blue Side) was used to make = spectrographic observations of QSO IR1321+055 in order to determine the = criteria by which the presence of BALs in the spectra may be predicted. = The observations were completed as planned, and no problems were = reported. 1.9 Completed WF/PC-2 6903 (WF/PC-2 Cycle 6 Calibration: = Decontamination) The WF/PC-2 was used to make several calibration bias = observations, internal flats, K-Spot images and dark calibration images = which are all part of a Cycle 6 decontamination procedure. No science = was interleaved with the dark exposures. The observations were completed = as scheduled, and no problems were reported. 1.10 Completed FOS/RD 6622 (UV Spectroscopy of DR1, the WO3 Star in = IC1613 and its surrounding Nebula) The Faint Object Spectrograph (Red Side) was used to make = ultraviolet spectrographic observations of the very luminous, WO3 = (Wolf-Rayet) class star DR-1 which is located in the metal-poor, dwarf, = irregular galaxy IC-1613. Observations were also be taken of its = surrounding nebula. The observations were completed as planned, and no = problems were reported. 1.11 Completed WF/PC-2 6909 (WF/PC-2 Cycle 6 Calibration: Earth Flats) The WF/PC-2 was used to make observations of the bright earth to = obtain streak flats which are used to construct high quality flat fields = for the WF/PC-2 filter set. The observations were executed as scheduled, = and no anomalies were noted. At this point, the safemade event described in 2.3 occurred, and the following scheduled proposals did not execute: 1. FOC/96 6349 (Snapshot search for High Redshift QSOs with Far- UV Flux) 2. FOS/BL/GHRS/1 6466 (Fundamental Properties of a DA-dM Detached Binary With Period=3D 3H 37M) 3. Three Sets of WF/PC-2 6253 (Pure Parallel Archive Program) 4. WF/PC-2 6362 (Supernova Host Galaxy) 5. FOS/BL 6465 (The UV Upturn of Super-Metal-Rich Globular Clusters) 6. FOS/BL 6539 (Ultraviolet Spectroscope of Seyfert 2 Nuclei) 7. Two Sets of WF/PC-2 6903 (WF/PC-2 Cycle 6 Calibration: Decontamination) 8. GHRS/2 6620 (Density Diagnostics for the Dynamic Atmospheres of LPV Stars) 9. WF/PC-2 6053 (Extreme UV Stars in Omega Cent) 10. GHRS/2 6534 (Be-Deficient Halo Stars: Implications For Galactic Chemical Evolution) 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 5 Successful : 5 Scheduled Reacquisitions : 13 Successful : 13 2.2 FHST Updates: Scheduled : 8 Successful : 8 2.3 Operations Notes: At 349/152802Z and as described in HSTAR 5845, the HST entered = Software Sumpoint Mode. As had occurred on October 30, the safemode test = for the RWA-1 momentum test failed, the wheel was commanded to off, and = safemode was entered. A RWA-1 power-on/torque test was successfully performed at = 349/2047Z. A Health and Safety load was uplinked and began executing at = 350/1700Z. Normal science operations are scheduled to resume at = 351/1645Z. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Nothing at this time. /DMH